OpenStudy (anonymous):

First one with correct answer will get fan and medal!!!!! Solve. 8 - 3x - x = 12 x = ?

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

cuz 8-3=5 5*3=15 15-3=12

OpenStudy (anonymous):

Can u help here too 3x – 6 = −2x + 9 x = ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

@angel12310 help?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

@ugh.math @shankar.smart can u help?

OpenStudy (anonymous):

wait its 3

OpenStudy (anonymous):

u sure

OpenStudy (anonymous):

3*3=9 9-6=3 -2*3=-6 -6+9=3

OpenStudy (anonymous):

Im sure

OpenStudy (anonymous):

ok one sec ill get another question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Josh has been running for 6 min and completes a lap in 7 min. Gabe has been running for 2 min and completes a lap in 9 min. How many more laps must Josh and Gabe complete to have run the same amount of time?

OpenStudy (anonymous):

Id think 63 mins. I might be wrong though

OpenStudy (anonymous):

because 63 is the LCM of both 7 and 9

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

i guess its 63 go with what u think

OpenStudy (anonymous):

@olivia.smiles @lordoftanks can u help?

OpenStudy (anonymous):

sure what can I help with :)

OpenStudy (anonymous):

@monsterfox1997 hey dude long time no see can u help? Josh has been running for 6 min and completes a lap in 7 min. Gabe has been running for 2 min and completes a lap in 9 min. How many more laps must Josh and Gabe complete to have run the same amount of time?

OpenStudy (anonymous):

I think its 63 because its LCM of both 7 and 9

OpenStudy (anonymous):

I agree also remember I you are pretty sure with an answer go with your first guess